# Knowledge in multi-robot systems: an interplay of dynamics, computation and communication

### v1

@@ Abstract
We show that the hybrid systems perspective of distributed multi-robot systems is compatible with logical models of knowledge already used in distributed computing, and demonstrate its usefulness by deriving sufficient epistemic conditions for exploration and gathering robot tasks to be solvable. We provide a separation of the physical and computational aspects of a robotic system, allowing us to decouple the problems related to each and directly use methods from control theory and distributed computing, fields that are traditionally distant in the literature. Finally, we demonstrate a novel approach for reasoning about the knowledge in multi-robot systems through a principled method of converting a switched hybrid dynamical system into a temporal-epistemic logic model, passing through an abstract state machine representation. This creates space for methods and results to be exchanged across the fields of control theory, distributed computing and temporal-epistemic logic, while reasoning about multi-robot systems.
@@ 1 Introduction
In this paper, we provide a connection between the hybrid systems approach to distributed multi-robot systems [34] and temporal epistemic logic via a variation of the runs and systems framework [18], which we construct by defining explicitly a state machine representation of both the system dynamics and the robot protocols. Bridging these fields enables us to combine well-known methods and results from control theory, distributed computing and temporal epistemic logic in a simple and harmonious fashion. This is a non-trivial connection not only because it connects a discrete and a continuous field of research; it also necessarily deviates from the established distributed computing principle that an optimal protocol always exists with respect to information exchange. We further demonstrate the usefulness of our approach by deriving the first ever sufficient epistemic conditions for solving the exploration, surveillance and approximate gathering robot tasks.
The distributed coordination of robots combines problems found in control theory [36, 26] with the problems commonly found in distributed computing [31]. While robotics is traditionally concerned with physical entities that must behave in accordance to their dynamical constraints, the distributed computing perspective arises once the robots must exchange information under varying difficulties associated to their communication capacities [1, 14]. For instance, understanding the impact of synchrony, coordination and message loss are essential for the efficient design and analysis of state-of-the-art multi-robot systems solving various tasks.
In [1] the authors famously showed that exact gathering is impossible by reducing the robot task to the consensus problem in distributed computing and deriving its well understood impossibility borders from there. In essence, this paper puts this reduction on a sound basis by deriving the distributed system implied by the control theoretic model and equipping it with the powerful tools found in epistemic logic.
Epistemic logic [24] is a powerful conceptual tool for reasoning about the uncertainty of agents in a system, both regarding facts and other agents’ epistemic states, making it especially suitable for the modeling of multi-agent systems, including distributed systems [23, 20, 29, 10, 21, 32]. A crucial feature of epistemic logic in multi-agent systems is that it provides a precise and formal representation of higher order epistemic attitudes, such as nested knowledge or nested belief, often crucial for the success of distributed tasks. For instance, coordinated actions are inevitably tied to nested knowledge, i.e., any agent needs to know to some extent what other agents know in order to successfully execute a coordinated behavior [32]. Another interesting feature of epistemic logic is that, when combined with temporal logic, it allows to express properties that must hold (and be known) throughout the execution, without focusing on the specific communication mechanisms leading to such knowledge. Thus, epistemic logic is a universal vehicle for expressing communication models, including those that rely on lack of explicit communication (also known as communication by silence) [20, 21].
@@
The control theoretic view on multi-agent systems has been developed by a number of authors, see e.g. the survey [4]. Similarly, logical views on hybrid systems have been considered, see e.g. [33]. As we mentioned also, logics, in particular epistemic logic [23] and temporal-epistemic logic [28] have been developed and applied to multi-agent systems, through in particular epistemic planning [17]. Our approach is the first one to seamlessly integrate three often complementary views: the ones from control theory, distributed systems and temporal epistemic logic.
1.1 Contributions
From the robot abstractions introduced, we define a representation of multi-robot system executions that is pertinent to both the control theory and distributed computing communities. From this common view, we derive a variant of the runs and systems framework wherein atomic propositions represent open sets in a topology, each corresponding to a region of a dense Euclidean space. We then use this epistemic model to provide the first epistemic analysis of key robot tasks, such as exploration, surveillance and approximate gathering, showing that the tasks are solvable under certain (epistemic) conditions. In particular, we prove that the sufficient conditions to solve approximate gathering mimic the sufficient conditions for the stabilizing agreement problem in distributed systems.
@@ 1.1 Contributions
This paper is, to the extent of our knowledge, the first framework that connects all three fields comprising control theory, distributed computing, and temporal epistemic logic, in a general setting that enables a comprehensive and integral approach to multi-robot systems.
1.2 Paper organization
@@
The paper is organized as follows. In Section 2 we introduce the basic concepts used for modeling multiple robots: first as a system of ordinary differential equations (Section 2.1), then as state machines (Section 2.2) and finally the correspondence of the two models is shown in Section 2.3. In Section 3, we derive an abstract distributed system from the robot models previously defined. Section 3.1 describes a scheduler that is compatible with both models, followed by matching definitions of system runs using state machines (Section 3.2) and switched systems (Section 3.3). Section 3.4 crucially connects the robotic abstraction to the epistemic model by turning the abstract robot state machine into a suitable epistemic frame. In Section 4, the epistemic frame is enriched into an epistemic model (Section 4.1) , which is then used for the epistemic analysis of robot tasks like simple exploration (Section 4.2), surveillance (Section 4.3) and approximate point gathering (Section 4.4). Finally, concluding thoughts are provided in Section 5.
@@ 2 The Robot Model
In this section, we introduce two multi-robot system models with distinct degrees of abstractions and subsequently show one model to be an abstraction of the other: the first is a description of robots as a hybrid dynamical system, close to the traditional approach within the control theory literature [25], the second is a state machine executing look-compute-move cycles, an abstraction commonly found in the distributed computing community [19]. This compatibility allows us to develop a common notion of scheduler for both the dynamical and the computational system perspectives, in Section 3, which will be used to define matching system runs.
@@ 2.1 Robots as Dynamical Systems
In the perspective of dynamical systems, robotic systems can modeled via differential equations that capture the evolution of their states over time. Hybrid systems are often used to distinguish the continuous state space of the physical aspects of a robot, from the discrete one of its computations. We first describe the model for a single robot, capable of movement and of computations.
@@
The ontic state corresponds to the physical state of the robot, while the epistemic state corresponds to the memory state of the robot. The distinction between the ontic and epistemic states allows for the physical and computational problems to be studied separately. As such, a robot’s behavior is given by a (timed) switched system [30], a particular case of hybrid system, as the control $\mathbf{u}[t]$ is possibly discontinuous over time.
Simplifying assumptions can be introduced so that one aspect, epistemic or ontic, is focused over the other. For instance, the epistemic evolution can be rewritten as an output-feedback controller so to recover the traditional control-theoretic point of view [26].
Example 2.2 (State feedback control).
Let a robot be a dynamical system with observer, $\dot{\mathbf{x}}=f(\mathbf{x},\mathbf{u})$ and $\mathbf{y}=g(\mathbf{x})$. An output feedback controller simulates the ontic evolution of the system $\mathbf{x}$ in order to obtain an estimation $\widehat{\mathbf{x}}$ of the state and regulate the input $\mathbf{u}$ of the original system.
$\dot{\widehat{\mathbf{x}}}$ $=f(\widehat{\mathbf{x}},\mathbf{u})$
$\mathbf{u}$ $=h(\widehat{\mathbf{x}},\mathbf{y})$
This controller corresponds to a robot where its epistemic state is an estimation of its ontic state, $\mathbf{e}(t)=\widehat{\mathbf{x}}(t)$, by following the same evolution, $\varphi=f$. The epistemic evolution uses the previous input $\mathbf{u}$ produced by itself for the simulation. The control accesses the observation $\mathbf{y}$, in order to compare with its own, i.e. $\mathbf{u}=h(\widehat{\mathbf{x}},\mathbf{y})=\mathbf{y}-g(\widehat{\mathbf{x}})$.
Alternatively, a point-mass dynamics could be imposed to the ontic evolution in order to obtain a system analogous to distributed computing. Moreover, by assuming a discrete-time dynamics one arrives at a computational model used in complexity theory, see for instance Sections 2.7 and 2.8 in [6].
Example 2.3 (Computing system).
Let a robot be a single computer, with a trivial (constant) ontic evolution $f(\mathbf{x}(t),\mathbf{u}[t])=0$ and an epistemic evolution representing the ’program’. In this context, the literature usually calls them agents.
$\dot{\mathbf{x}}(t)$ $=f(\mathbf{x}(t),\mathbf{u}[t])=0$
$\mathbf{y}(t)$ $=g(\mathbf{x}(t))$
$\mathbf{e}[t+1]$ $=\varphi(\mathbf{e}[t],\mathbf{y}[t])$
$\mathbf{u}[t]$ $=h(\mathbf{e}[t],[t])$
As $\mathbf{x}(t)$ is some constant, $\mathbf{y}(t)$ is also constant. $\varphi$ now only depends on $\mathbf{e}[t]$ and can be seen as the memory of the single computer. The initial external input can be provided via the constant value of $\mathbf{y}(t)$.
Remark 2.4 (From continuous to discrete).
Due to the physical limitations in memory of a physical computing device, we assume that the observation function $g$ has a finite domain $\mathcal{D}(g)$, which corresponds to the digital representation of the continuous quantities it may be capable of measuring. This captures the role of sensors in robotics, such as LIDARs that discretize the environment in distance measurements or temperature sensors that ultimately convert their values to binary via an ADC measuring the voltage drop over a heat sensitive element. The size of $\mathcal{D}(g)$ will depend on how the robots measure their surroundings and derive computable information from it, but we assume that $g$ spans the whole set $\mathcal{D}(g)$, i.e., it is surjective.
A finite and discrete abstraction will be useful in Section 4, where the propositional modal language does not provide quantifiers, restricting to the representation of finite sets of facts. Note that our goal is to derive formulas depending on known facts about the world, therefore reasoning about the robots’ knowledge about facts representing the ontic world in their local memories.
@@
We now consider the modifications required for modeling a multi-robot system where interactions increase in complexity with the added notions of communication and perception among robots. Note that the distinction between communication and perception is a consequence of the distinction between ontic and epistemic states, as both are required to represent the possible information to be exchanged.
Definition 2.5 (Communication and perception of robots).
Let $\mathbf{r}$ and $\mathbf{r}^{\prime}$ be robots. The perception function $\eta$ provides to robot $\mathbf{r}$ an estimation $\widehat{\mathbf{x}}_{\mathbf{r}\mathbf{r}^{\prime}}$ of robot $\mathbf{r}^{\prime}$’s ontic state, and the communication function $\lambda$ provides robot $\mathbf{r}$ with an estimation $\widehat{\mathbf{e}}_{\mathbf{r}\mathbf{r}^{\prime}}$ of the epistemic state of robot $\mathbf{r}^{\prime}$.
@@
Communication and perception complement the observation step with information about the other robots, which in turn is used by the epistemic evolution.
$\begin{split}\mathbf{y}_{\mathbf{r}}(t)=g(\mathbf{x}(t),\{\eta_{\mathbf{r}}(\mathbf{x}_{\mathbf{r}^{\prime}}(t))\},\{\lambda_{\mathbf{r}}(\mathbf{e}_{\mathbf{r}^{\prime}}[t])\})\hskip 30.00005pt\text{(multi-robot observation)}\end{split}$ (7)
Note that Definition 2.5 places communication and perception as two types of interaction that are processed together by the epistemic evolution of the robot in Equation 3. This understanding will match the behavior of observation operations in the perspective of robots as computing systems.
@@ 2.2 Robots as Mobile Computing Systems
By discerning the physical dynamics of the robot from the dynamics of its computations, it becomes evident that the epistemic evolution matches the role of an algorithm, where the epistemic state is the memory being operated on. Now we will explore how the robots described in Definition 2.1 can be abstracted as mobile computing systems executing look-compute-move cycles in a state machine.
Definition 2.6 (look-compute-move cycle [19]).
The look-compute-move model decomposes the behavior of robots in an environment as cycles that execute three active steps and one inactive step. The environment consists of the universe in which the robots move (for example a graph or a continuous region of space), and some adversarial behavior influencing the robots’ actions. The adversarial definition includes at least a scheduler deciding on which robot takes the next step in the look-compute-move cycle.
@@
    Look. The robot makes use of it sensors to retrieve information from the environment. A snapshot displays the physical state of the other robots, possibly limited by a visibility radius.
@@
    Compute. The information obtained in the Look step is used to locally compute the succeeding memory state. All robots execute the same deterministic algorithm.
@@
    Move. The robot changes its physical state (based on its current memory state), which most commonly consists of its position in the environment.
@@
Uncertainty in the behavior of each step, such as the accuracy of the information or the actual moved distance, are possible choices for adversarial behavior in the LCM model. Some variations are depicted in Appendix A.
Note that the physical dynamics of a robot are only relevant when interacting with the environment, by updating (move) or by retrieving (look) its state. The computational aspects of a robot are encapsulated by the compute step, that matches the processing of the acquired information and the generation of commands to be executed.
The simplest models of robots running lcm cycles are not capable of interaction besides observing the existence and physical state of each other via look operations. Therefore, we will consider the enhanced model of luminous robots, equipped with programmable lights for communication and memory. Note that this is a parallel requirement to the communication and perception introduced in Definition 2.5.
Definition 2.7 ($\mathcal{LUMI}$ model [12]).
The LUMInous robot model consists of a group of homogeneous robots operating lcm cycles. They are modeled as points in an environment $\mathcal{E}$ that operate without central control. Each robot possesses a set of external lights that can be modified in order to express some information for communication, as they can be observed by other robots, or for memory, as they are preserved throughout rounds. The physical state, including the lights, are updated during a move step. During look operations, each robot obtains a snapshot of the environment $\mathcal{E}$, which provides the physical states, including the lights, of all other robots. In the weakest version, robots with the same lights are indistinguishable and have no identity. They are disoriented and possess no global reference frame.
Remark 2.8.
Note that the communication made available through the $\mathcal{LUMI}$ model is indirect, as any message provided by a robot must be actively observed by another one for it to be received. Here, perception and communication are combined in a single interaction. The outgoing communication is entirely comprised by the actions that the robot can make, i.e., set the lights according to its new state. Symmetrically, the incoming communication is entirely described by the observations made by a robot.
Now, we describe the state machine executed by a robot operating under the $\mathcal{LUMI}$ model. We consider that a robot $\mathbf{r}$ exists in an environment $\mathcal{E}$ and executes $\mathcal{A}$ during the compute step of the lcm cycle. The evolution of $\mathcal{E}$ comprehends both of the move and look steps, where the ontic updates of all robots are executed and the possible observations retrieved.
@@
$\mathcal{A}=\langle S_{\mathbf{e}_{\mathbf{r}}},S_{\mathbf{y}_{\mathbf{r}}},S_{\mathbf{u}_{\mathbf{r}}},g_{\mathbf{r}},\varphi_{\mathbf{r}},h_{\mathbf{r}}\rangle$
where
@@
The environment evolves according to the state machine
$\mathcal{E}=\langle S_{\mathcal{E}},I_{\mathcal{B}},\{S_{\mathbf{u}_{\mathbf{r}}}\}_{\mathbf{r}\in\Pi},\{S_{\mathbf{y}_{\mathbf{r}}}\}_{\mathbf{r}\in\Pi},F_{\mathcal{E}},G_{\mathcal{E}}\rangle$
where
@@
This depiction allows us to focus on the computations executed locally, while the environment abstracts all ontic states and updates into a single entity, with which all robots’ state machines interact. The look and move are split into distinct functions that interface the environment and robot: one of preparation of the data ($h_{\mathbf{r}}$ and $G_{\mathcal{E}}$), and one of usage of the data, $F_{\mathcal{E}}$ and $g_{\mathbf{r}}$. This split will facilitate stating the correspondence between computational and dynamical perspectives in Section 2.3.
This abstraction will provide the algorithmic perspective of a robotic system essential to obtaining a principled representation of the knowledge available in Section 4.
2.3 Correspondence of the Dynamical and Computational Perspectives
We now have the language to state the compatibility of the dynamical robot model described in Section 2.1 and the computational model described in Section 2.2. Let $\mathbf{r}$ be a robot modeled as a hybrid system according to Definition 2.1 with communication and perception as in Definition 2.5. Let $\mathbf{r}^{\prime}$ be a luminous robot, as in Definition 2.7 executing lcm cycles as in Definition 2.9, and let $\mathcal{E}$ be the environment in which $\mathbf{r}^{\prime}$ executes, as in Definition 2.10. We have that:
Proposition 2.11 (Correspondence of discretized state spaces).
The following functions that map the space of $\mathbf{r}$ to that of $\mathbf{r}^{\prime}$ are surjective.
  1. (i)
    $\Phi_{\mathbf{e}}:\mathbf{e}_{\mathbf{r}}\to S_{\mathbf{e}_{\mathbf{r}^{\prime}}}$, from epistemic states to memory states;
  2. (ii)
    $\Phi_{\mathbf{y}}:\mathbf{y}_{\mathbf{r}}\to S_{\mathbf{y}_{\mathbf{r}^{\prime}}}$, from observations to discrete observations, as per Remark 2.4;
  3. (iii)
    $\Phi_{\mathbf{u}}:\mathbf{u}_{\mathbf{r}}\to S_{\mathbf{u}_{\mathbf{r}^{\prime}}}$, from a discrete control to discrete actions;
  4. (iv)
    $\Phi_{\mathcal{E}}:\mathbf{x}_{\mathbf{r}}\to S_{\mathcal{E}_{\mathbf{r}^{\prime}}}$, from ontic states to environment states.
Proposition 2.12 (Correspondence of dynamics).
For any execution of $\mathbf{r}$ over a time interval $[t,t+\Delta]$ corresponding to one lcm cycle of $\mathbf{r}^{\prime}$, we have that:
  1. (i)
    $\varphi_{\mathbf{r}}:\mathbf{e}_{\mathbf{r}}[t]\times\mathbf{y}_{\mathbf{r}}[t]\to\mathbf{e}_{\mathbf{r}}[t+1]$ corresponds to $\varphi_{\mathbf{r}^{\prime}}:S_{\mathbf{e}_{\mathbf{r}^{\prime}}}\times S_{\mathbf{y}_{\mathbf{r}^{\prime}}}\to S_{\mathbf{e}_{\mathbf{r}^{\prime}}}$. Also, $h:\mathbf{e}_{\mathbf{r}}[t]\to\mathbf{u}_{\mathbf{r}}[t]$ corresponds to $h_{\mathbf{r}^{\prime}}:S_{\mathbf{e}_{\mathbf{r}^{\prime}}\to S_{\mathbf{u}_{\mathbf{r}^{\prime}}}}$. That is, the epistemic and control functions match the robot update.
  2. (ii)
    $f_{\mathbf{r}}(\cdot)$ corresponds to $F_{\mathcal{E}}(\cdot)$. That is, the ontic evolution corresponds to the environment update.
  3. (iii)
    $g_{\mathbf{r}}(\cdot)$ corresponds to $g_{\mathbf{r}^{\prime}}(G_{\mathcal{E}}(\cdot))$. That is, the observation functions match.
Proposition 2.13 (Correspondence of interactions).
The communication $\lambda_{\mathbf{r}_{1}\mathbf{r}_{2}}(\cdot)$ and perception $\eta_{\mathbf{r}_{1}\mathbf{r}_{2}}(\cdot)$ functions, in Equation 7 of Definition 2.5, alongside the general environment perception $g_{\mathbf{r}}$ correspond to the components of the function $G_{\mathcal{E}}(\cdot)$ generating observations from the environment in Definition 2.10. That is, the communication and perception functions correspond to the observation generation in the environment state machine.
3 Distributed Multi-Robot Systems and Epistemic Frames
In this section, we will define the construction of system runs, and support it by a general notion of scheduler, applicable to both the dynamical and computational perspectives provided in Section 2. A general epistemic frame for multi-robot tasks will then be introduced using the system runs as its building block. The epistemic frames will be later specialized, in Section 4, to runs and systems models fitting different robot tasks.

### v2

@@ Abstract
In this paper, we provide a framework integrating distributed multi-robot systems and temporal epistemic logic. We show that continuous-discrete hybrid systems are compatible with logical models of knowledge already used in distributed computing, and demonstrate its usefulness by deriving sufficient epistemic conditions for exploration and gathering robot tasks to be solvable. We provide a separation of the physical and computational aspects of a robotic system, allowing us to decouple the problems related to each and directly use methods from control theory and distributed computing, fields that are traditionally distant in the literature. Finally, we demonstrate a novel approach for reasoning about the knowledge in multi-robot systems through a principled method of converting a switched hybrid dynamical system into a temporal-epistemic logic model, passing through an abstract state machine representation. This creates space for methods and results to be exchanged across the fields of control theory, distributed computing and temporal-epistemic logic, while reasoning about multi-robot systems.
@@ 1 Introduction
We formulate in this paper a framework connecting a hybrid systems approach to distributed multi-robot systems and temporal epistemic logic. This forms a bridge whose value is twofold: we bring the success of logical analysis of tasks in distributed computing [17] to distributed robotics, and we generalize the multi-robot model from state machines to dynamical systems [33], drawing a much closer frontier to real-life applications.
The distributed coordination of robots combines problems found in control theory [35, 25] with the problems commonly found in distributed computing [30]. By clearly separating the roles of the physical and computational parts of a robot, we are capable of showing a translation layer from distributed computing and control theory. While robotics is traditionally concerned with physical entities that must behave in accordance to their dynamical constraints, the distributed computing perspective arises once the robots must exchange information under varying difficulties associated to their communication capacities [1, 12]. For instance, understanding the impact of synchrony, coordination and message loss are essential for the efficient design and analysis of state-of-the-art multi-robot systems solving various tasks.
Our modeling is extended for the multi-robot scenario, where we introduce the notion of a scheduler as a path in the space of executions, so to tie local to global time. Bridging these fields enables us to combine well-known methods and results from control theory, distributed computing and temporal epistemic logic in a simple and harmonious fashion. Our framework successfully captures the spatial information used to reason about multi-robot tasks, namely exploration and gathering, as it is constructed with the memory of the robots during their computations. We devise a logical model consisting of a variation of the runs and systems framework, which we construct by defining explicitly a distributed state machine for both the system dynamics and the robot protocols. We further demonstrate the usefulness of our approach by deriving the first ever sufficient epistemic conditions for solving the exploration, surveillance and approximate gathering robot tasks.
In [1] the authors famously showed that exact gathering is impossible by reducing the robot task to the consensus problem in distributed computing and deriving its well understood impossibility borders from there. In essence, the present paper puts this reduction on a sound footing by deriving the distributed system implied by the control theoretic model and equipping it with the tools found in epistemic logic.
@@
The control theoretic view on multi-agent systems has been developed by a number of authors, see e.g. the survey [4]. Similarly, logical views on hybrid systems have been considered, see e.g. [32]. As we mentioned also, logics, in particular epistemic logic [22] and temporal-epistemic logic [27] have been developed and applied to multi-agent systems, through in particular epistemic planning [15].
Epistemic logic [23] is a powerful conceptual tool for reasoning about the uncertainty of agents in a system, both regarding facts and other agents’ epistemic states, making it especially suitable for the modeling of multi-agent systems, including distributed systems. It was adapted for and introduced to distributed systems in [22], subsequently employed to characterize: the firing rebels with relay in [19], two agent approximate consensus in [28], sufficient conditions for stabilizing consensus in [8], optimal consensus with respect to message transmission in[20] and different notions of knowledge in [31], to name just a few. A crucial feature of epistemic logic in multi-agent systems is that it provides a precise and formal representation of higher order epistemic attitudes, such as nested and/or common knowledge, often crucial for the success of distributed tasks. For instance, coordinated actions are inevitably tied to nested knowledge, i.e., any agent needs to know to some extent what other agents know in order to successfully execute a coordinated behavior [31]. Another interesting feature of epistemic logic is that, when combined with temporal logic, it allows to express properties that must hold (and be known) throughout the execution, without focusing on the specific communication mechanisms leading to such knowledge. Thus, epistemic logic is a universal vehicle for expressing communication models, including those that rely on lack of explicit communication (also known as communication by silence) [19, 20]. Our approach is the first one to seamlessly integrate these three complementary views: the ones from control theory, distributed systems and temporal epistemic logic.
@@
Contributions. We present, to the extent of our knowledge, the first interdisciplinary framework that connects all three fields of control theory, distributed computing, and temporal epistemic logic, in a general setting that enables a comprehensive and integral approach to multi-robot systems. We showcase the potential of our framework by defining a simple yet sufficiently expressive temporal epistemic language that allows us to express multi-layered knowledge of space. In particular, this enables a precise formalization for the completion of spatial robotic tasks, such as exploration and approximate gathering. More precisely, our main contributions are:
  • We define a robot abstraction representing multi-robot system (Definition˜2.3) executions compatible with to both the control theory (Definition˜2.1) and distributed computing (Definitions˜2.4 and 2.5) frameworks.
  • We provide a temporal epistemic logic framework (Definition˜4.1), based on the runs and systems framework, for expressing robot exploration as knowledge of atomic propositions that represent open sets of a compact topological space.
  • We provide an epistemic analysis of exploration (Definition˜4.4), and approximate gathering (Definition˜4.34). Furthermore, we characterize exploration with cooperative termination (Theorem˜4.14), and we prove that the sufficient conditions to solve approximate gathering mimic the sufficient conditions for the stabilizing agreement problem in distributed systems (Theorem˜4.36).
@@
Paper Organization. In Section˜2 we introduce the basic concepts used for modeling multiple robots: first as a system of ordinary differential equations (Section˜2.1), then as state machines (Section˜2.2). In Section˜3, we introduce time paths as a geometric model relating global and local times (Section˜3.1), followed by their usage in the definitions of runs in both dynamical (Section˜3.2) and computational models (Section˜3.3). We conclude by showing that the computational content of these runs match (Section˜3.4). Section˜4, the main contribution of the paper, connects the robotic abstraction to the epistemic modeling by turning the abstract robot state machine into a suitable epistemic frame. The latter is enriched into an epistemic model (Section˜4.1), which is then used for the epistemic analysis of robot tasks such as simple exploration (Section˜4.2) and approximate gathering (Section˜4.4). Finally, concluding thoughts are provided in Section˜5.
@@ 2 The Robot Model
In this section, we introduce two multi-robot system models with distinct degrees of abstractions, and subsequently show one model to be an abstraction of the other. The first is a description of robots as a switched hybrid dynamical system, close to models found in the control theory literature [24]. The second is a state machine executing look-compute-move cycles, an abstraction commonly found in the distributed computing community [18].
@@ 2.1 Robots as Dynamical Systems
From the dynamical systems perspective, a robotic system can modeled via differential equations that describe the evolution of its state over time. Hybrid systems combine a continuous state space, associated to the physical features of a robot, with a discrete space, more natural to capture computations. We start by describing the model for a single robot.
@@
The ontic state corresponds to the physical state of the robot, while the epistemic state corresponds to the memory state of the robot. The distinction between the ontic and epistemic states allows for physical and computational problems to be described separately. As such, a robot’s behavior is given by a (hybrid) switched system [29], a particular case of hybrid system, as the control $\mathbf{u}[t]$ is possibly discontinuous over time.
Note that Definition˜2.1 also encompasses more traditional models of robots in the control-theoretic literature, such as output-feedback controllers [25]. This can be achieved by using the epistemic evolution $\varphi$ as a simulator that imitates the dynamics of the model.
Remark 2.2 (Discrete Observations).
Due to the physical limitations in memory of any physical computing device, we assume that the observation function $g$ has a finite domain $\mathcal{D}(g)$, which corresponds to the digital representation of the continuous quantities it is capable of measuring.
@@
We now consider the modifications required for modeling a multi-robot system where interactions increase in complexity with the added notions of communication and perception among robots. Note that the distinction between communication and perception is a consequence of the distinction between ontic and epistemic states, as both are required to represent the possible information to be exchanged. For two robots $\mathbf{r}$ and $\mathbf{r}^{\prime}$, they are defined via a pair of functions.
@@
The perception is $\widehat{\mathbf{x}}_{\mathbf{r},\mathbf{r}^{\prime}}$ is the estimation of the ontic state of robot $\mathbf{r}^{\prime}$ by robot $\mathbf{r}$. Similarly, The communication function generates the estimation $\widehat{\mathbf{e}}_{\mathbf{r},\mathbf{r}^{\prime}}$ of the epistemic state of robot $\mathbf{r}^{\prime}$ by robot $\mathbf{r}$. Moreover, the introduction of multiple robots operating in the system brings forward the notions of "local" and "global" execution time, which will be formalized in Section˜3.1.
Communication and perception complement the observation step with information about the other robots. These multiple pieces of information are later processed together by the robot’s epistemic evolution, matching the intuition of observations as simply inputs to a computing system. This leads to the definition of a hybrid multi-robot system.
Definition 2.3 (Hybrid multi-robot system).
A hybrid multi-robot system $\mathcal{H}$ consists of $N$ robots, capable of observing the environment and interacting with the other robots through perception and communication. For a given sequence of activations $p$, its evolution is given by:
$\mathbf{U}[t]$ $=H(\mathbf{E}[t],t)$ (global control) (10)
This definition is a consequence of the evolution of single robots alongside the interaction functions, as in Definitions˜2.1, 5 and 6. We now consider a generic scheduler that provides the set of active robots at each time $t$, defining the modes of the switched system, and that will be made precise with the introduction of time paths in Section˜3.1.
@@ 2.2 Robots as Mobile Computing Systems
We provide now a state machine abstraction that allows us to focus on the computational component of the robots in Definition˜2.1. Robots are algorithmically described as mobile entities that follow a cycle of look-compute-move operations as they interact with the environment. This is known as the LCM cycle[18]:
@@
    look. The robot retrieves a snapshot of the ontic state of other robots in the environment.
@@
    compute. The robot computes its next epistemic state from the information in look.
@@
    move. The robot updates its physical state based on its current epistemic state.
@@
All robots operate in the same environment, e.g. a graph or continuous region of space. Uncertainty in the behavior of each step, such as the accuracy of the information or the actual moved distance, are possible choices for adversarial behavior in the LCM model. Some variations are depicted in Appendix A.
Two models for robots running LCM cycles are predominant in the literature: oblivious, introduced in [36], and luminous, introduced in [10]. They contrast in the means for communication and memory made available to the robots. Robots under the oblivious robot model, $\mathcal{OBLOT}$, have no memory and can only obtain information from observing other robots in the look step. Robots under the luminous robot model, $\mathcal{LUMI}$, are enhanced with a set of programmable lights that can be user for both (indirect) communication and storing information for future compute steps. The lights are a physical element representing the exposed memory of a robot, and hence can be observed with the rest of the ontic state of a robot during look operations and are updated during a move operation. We will use the $\mathcal{LUMI}$ model, as we require the lights to handle both communication and perception. We instantiate the robot $\mathbf{r}$ in an environment $\mathcal{E}$ executing $\mathcal{A}$ as follows.
@@
$\mathcal{A}=\langle S_{\mathbf{e}_{\mathbf{r}}},S_{\mathbf{y}_{\mathbf{r}}},S_{\mathbf{u}_{\mathbf{r}}},g_{\mathbf{r}},\varphi_{\mathbf{r}},h_{\mathbf{r}}\rangle$ where
@@
The environment executes the state machine $\mathcal{E}=\langle S_{\mathcal{E}},I_{\mathcal{B}},\{S_{\mathbf{u}_{\mathbf{r}}}\}_{\mathbf{r}\in\Pi},\{S_{\mathbf{y}_{\mathbf{r}}}\}_{\mathbf{r}\in\Pi},F_{\mathcal{E}},G_{\mathcal{E}}\rangle$ where
@@
Definition˜2.4 lets us focus on the computations executed locally, while Definition˜2.5 abstracts the ontic states and updates of all robots in a single machine interacting with the individual computations. look and move are split into distinct functions at the interface between the environment and each robot: the data is “prepared” by $h_{\mathbf{r}}$ and $G_{\mathcal{E}}$, and “consumed” by $F_{\mathcal{E}}$ and $g_{\mathbf{r}}$.
A computational multi-robot system $\mathcal{C}$ is then a set of robots and one environment state machine, i.e., $\mathcal{C}=(\{\mathcal{A}_{\mathbf{r}}\}_{\mathbf{r}\in\Pi},\mathcal{E})$. This separation will facilitate stating the correspondence between computational and dynamical perspectives in Section˜3.4.
3 Schedulers and System Runs
In this section, we define a scheduler (Section˜3.1) that dictates the possible activations of the distributed system. System runs are thus introduced, in both the dynamical (Section˜3.2) and computational (Section˜3.3) definitions. The runs will be later used as building blocks to generate a general epistemic frame for multi-robot tasks in Section˜4.